OpenStudy (anonymous):

Find the solution set of this inequality. Enter your answer in interval notation using grouping symbols. |-2x+1|<=5

OpenStudy (anonymous):

-2x+1<5 -2x<4 x>2 (2,int) 2x-1<5 2x<6 (3,inf)

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

wait NO [2,inf) and (-inf,3]

OpenStudy (anonymous):

|x|=+x and also -x so you have to solve two systems

OpenStudy (anonymous):

i cant put letters tho

OpenStudy (anonymous):

\[|-2x+1|\le5\] First, we remove the absolute value sign by rewriting the equation as follows: \[-2x+1 \le \pm5\] Now we can solve. First solution: \[-2x \le 4\] x > -2 Second solution: \[-2x \le -5\] x > 5/2 Hope this helps: http://www.tutorsean.net

OpenStudy (anonymous):

Oops. On the second one, it is x > 3.